Status Report

NASA Space Station On-Orbit Status 25 Jun 2003

By SpaceRef Editor
June 25, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 25 Jun 2003

All ISS systems continue to function nominally, except as noted previously or below.

CDR Yuri Malenchenko completed a major IFM (in-flight maintenance) to restore the functionality of the VFS-3M video-photometric system of the Russian Molniya-SM (GFI-10) geophysical experiment payload.  The IFM was supported by tagups with a ground specialist.  [Major tasks of the activity were replacement of the battery on the VFS-3M processor board, replacing the hard drive and resetting the BIOS parameters in the electronics module, then testing the new disk drive via the Agate display monitor, with optional replacement of the Fujitsu magneto-optical drive if found necessary.  In January this year, a test of the VFS-3M twin-lens video-photometric system had revealed a failure of the computer subsystem of the VFS electronics module (ME).  The  task today was a second attempt to restore the system, after Nikolai Budarin had already worked on it on 3/31.  The Molniya VFS-3M is used for studying atmospheric, ionospheric and magnetospheric electromagnetic interaction related to storms and seismic activities.]

Malenchenko deferred planned work on the hardware upgrade of the BRPK condensate separation and pumping units.  The plan is to replace the old pipe conduits, including safety valves, with new units.  [The latter have no filters and use higher safety valve trip pressures.  Purpose of the upgrade is to prevent unwanted flow of the fluid through the bypass line during condensate transfer cycles.]

The crew performed the second part of this week’s maintenance work on the RED (resistive exercise device) to swap out the Flexpacks of the two RED canisters with new units flown up on 11P.  Today’s 3-hr. IFM operation dealt with canister #1010.  [Can #1009 had received its new Flexpack on 6/23.  The replaced Flexpack units will be inspected to ascertain how many good packs are available for use as spares when the current Flexpacks fail.  The inspection is scheduled as a task list item between today and the weekend, and will be hard-scheduled during Week 9.  The replaced units from 1009 and 1010 will be used in canisters 1001 and 1003 in a separate IFM once these cans are considered failed.]

FE/SO Ed Lu also conducted the periodic (every other week) inspection of the RED and monthly bolt tightening (as required).

Dr. Lu collected and stowed the two passive FMK (formaldehyde monitoring kit) sampling badges which he had deployed in the Lab and SM on 6/23.

Then, he unstowed and installed the equipment for the U.S. PHS (periodic health status) with blood labs exam, scheduled for tomorrow.  Yuri similarly set up the test equipment for tomorrow’s periodic Russian MO-10 “Hematokrit” testing.  [While PCBA analyzes total blood composition, MO-10 particularly measures the hematocrit (red blood cell mass) value of the blood (as a well-known phenomenon of space flight, red blood cell mass {normal range: 30-45%} tends to go down over time).

Ed also conducted the daily routine maintenance of the SM’s SOZh environment control & life support system and prepared the daily “delta” file for updating the inventory management system (IMS).

Yuri completed his regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) payload which experiments with growth and development of plants under spaceflight conditions in the Lada-2 greenhouse.

At 12:25pm EDT, the crew successfully conducted an interactive audio/video PAO ship-to-ship call with the crew of the “Aquarius” underwater research facility of the NEEMO 5 (NASA Extreme Environment Mission Operations) mission from June 16-29.  [Aquarius, similar in size to the ISS Service Module (SM) and owned by NOAA, is the only undersea research laboratory in the world.  The 45-ft long by 13-ft wide habitat sits on the ocean floor three miles off Key Largo, Florida, near deep coral reefs, 62 ft beneath the surface.  The four-member NASA crew is headed by former ISS astronaut Dr. Peggy Whitson who, having spent 185 days in space, is well experienced to compare life in space with life under water.  The other crewmembers are Clay Anderson, Dr. Garret Reisman and Dr. Emma Hwang.]

The Science Officer was thanked for his excellent job with yesterday’s completely successful Ultrasound activity.  [The ultrasound video quality was sufficient for effective remote guidance and image capture for both clinical capabilities and potential research purposes.  This first demonstration of the ability to conduct remote guidance to/from the JSC Telescience Support Center (TSC) has great potential for future HRF (human research facility) research experiments, and also as a backup resource for Space Medicine objectives.]

The crew completed their regular daily 2.5-h program of physical exercise, on TVIS treadmill, RED expander and, for Yuri, on the Russian VELO cycle ergometer with load trainer.

Ed Lu reported that his heart rate monitor (HRM) battery has failed and that he has donned a new HRM, sent up on 11P/Progress.  The old HRM will be returned to Houston for refurbishment.  [The HRMs are pulse-sensitive wrist watches worn by the crew to monitor heart rate and control exercise level during daily workout.  They are the same as those flown on the Shuttle.  The accurately measured heart rate is transmitted to receivers located on the TVIS treadmill, CEVIS ergometer, etc.]

Moscow continues to study further troubleshooting of the Klest-140ST-M TV camera mounted externally on the SM aft end, pointing rearward (+X direction for SM).  [This camera is used for rendezvous and docking with the SM aft port.  Impacts of a failure to be determined.]  

Today’s CEO (crew earth observation) targets, limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, and including the targets of the Lewis & Clark 200-year memorial locations, were Barcelona, Spain (nadir pass), El Paso, Texas (nadir pass), Weeks Creek, California (landslide research site: to document this small site the ground suggested a detailed nadir mapping swath from the coast inland as far as San Francisco Bay), Guangzhou, China (panoramas of this metropolitan region situated on the great estuary of the Pearl River), Manila, Philippines (regional views left of track of Pinatubo volcano and coastline environment of Manila), Himalaya Mts (Dynamic event.  Cloud free conditions for this pass along the western sector of the Himalaya chain), Singapore (synoptic view of the southern tip of the Malay Peninsula, Strait of Malacca and Sumatraright of track), Eastern Mediterranean haze (somewhat reduced visibility reported along track at Venice, Athens and Beirut airports.  Crew was advised to shoot left and right of track for any detail in fair-weather photochemical smog accumulations), Tel Aviv, Israel (panoramas of the Levant, left and right of track), and Tucson, Arizona (oblique view [looking right] of this major fire that has burned 250,000 acres and 200 homes.  The blaze is slowly being brought under control on the steep slopes of the Catalina Mts).
CEO images can be viewed at the websites

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:48pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (18 amp mode).  Vozdukh CO2 scrubber is On (in Manual Mode 3/5).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 28.7; ppO2 (mmHg) — data invalid; ppCO2 (mmHg) — data invalid.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.9.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 22.0.
  • Node:  Pressure (mmHg) — 745.80; temperature (deg C) — 24.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 747.44; temperature (deg C) — 23.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 747.54; temperature (deg C) — 30.4; shell heater temp (deg C) — 25.3, ppO2 (mmHg) — 172.3; ppCO2 (mmHg) — 4.2.
  • PMA-1:  Shell heater temp (deg C) — 25.0.
  • PMA-2:  Shell heater temp (deg C) — 14.8.

(n/a = data not available)

Propulsion System (PS):

  • Total propellant load available: 3803 kg (8384 lb) as of 6/19  [SM(774) + FGB(2447) + Progress M(182) +Progress M-1(400)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in AutoTrack.
  • SM batteries:  Battery #4 is disconnected (failed 6/16); all other batteries (7) are in “Partial Charge” mode.
  • FGB batteries:  Battery #1 is in “Cycle” mode; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-1 is operating; INT-2 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is Off; PL-1 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).Management.
  • Solar Beta angle:  -23.7 deg (magnitude decreasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally [IAC-1 (internal audio controller #1) being analyzed after self-test error.  IAC-2 is prime.]
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:32am EDT [= epoch]):

  • Mean altitude — 388.0 km
  • Apogee — 392.7 km
  • Perigee — 383.2km
  • Period — 92.31min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0007015
  • Orbits per 24-hr. day — 15.60
  • Solar Beta Angle — -23.7 deg (magnitude decreasing)
  • Mean altitude loss in last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 26229
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.